OpenStudy (anonymous):

a ball is dropped from a cliff that is 224 feet high. the distance S (in feet) that it falls in t seconds is given by the formula S=224-16t^2. how many seconds (to tenths) will it take for the ball to hit the ground

OpenStudy (zzr0ck3r):

-26t^2+224=0 26t^2 = 224 t^2=224/26 t = sqrt(224/26)

OpenStudy (zzr0ck3r):

understand?

OpenStudy (anonymous):

Much better, thank you!